# An Interdisciplinary and Cross-Task Review on Missing Data Imputation

### v1

@@ 6.2 Link Prediction and Graph Completion
Graph completion, which is more general than link prediction, aims to enrich the entire graph by adding not only edges but also missing nodes, attributes, or labels so that the resulting graph is more complete and useful for downstream tasks. A general graph is formulated as $G=(V,E,\mathbf{X})$, where $\mathbf{X}$ denotes the matrix of node attributes. Graph completion may recover the missing edges as well as the missing values in $\mathbf{X}$, or predict where a new node should be inserted into the graph (Wu et al., 2023). Within the tasks of graph completion, knowledge graph completion (KGC) is best known. KGC is the task of predicting missing facts in a knowledge graph, typically by inferring new relationships between entities to answer queries like (subject, predicate, ?) or (?, predicate, object). There have been many review papers on KGC (Chen et al., 2020; Shen et al., 2022; Zamini et al., 2022; Cai et al., 2022a; Liang et al., 2023; Li et al., 2025a). In (Li et al., 2025a), KGC methods are organized into embedding-based methods, path-based methods, neural network-based methods, and LLM-based methods (Wang et al., 2022a; Wei et al., 2024; Xu et al., 2024a; Li et al., 2024a; Zhang et al., 2024b; Yao et al., 2025).
@@ 6.3 Time Series Imputation
The task of time series imputation is to recover the missing values in a single multivariate time series (e.g., a matrix $\mathbf{X}\in\mathbb{R}^{d\times T}$) or a union of independent multivariate time series (e.g., a union of matrices $\mathbf{X}_{i}\in\mathbb{R}^{d\times T_{i}}$, $i=1,\ldots,K$). Compared to the standard missing data imputation, time series data imputation should consider the auto-correlation, trend/seasonality, and lagged cross-correlation. Time-series imputation keeps critical real-time systems (power grids, ICU monitors, high-frequency trading) from blinding when sensors drop out, preventing cascading failures or false alarms. It also unlocks the full value of expensive historical records—satellite climate data, industrial IoT (internet of things) logs, retail POS streams—so forecasting, digital twins and compliance audits can run on complete, audit-ready data instead of discarding gap-ridden sequences. Recent advances in time series imputation are based on recurrent neural networks (Cao et al., 2018) and diffusion models (Tashiro et al., 2021). There have been a few review papers on time series imputation (Pratama et al., 2016; Weerakody et al., 2021; Ribeiro and de Castro, 2022; Zainuddin et al., 2022; Fang and Wang, 2020; Wang et al., 2024a; Qian et al., 2025). In the past ten years, deep learning methods, especially deep generative models, have shown impressive performance in time-series imputation. Fang and Wang (2020) wrote a survey paper of deep learning-based imputation methods for time series (Luo et al., 2018; Gupta and Beheshti, 2020; Park et al., 2023). Kazijevs and Samad (2023) focused on time series health data and provided a review with numerical comparisons for a few deep learning imputation methods. Jin et al. (2024) reviewed graph neural network-based methods for time series imputation. Du et al. (2024b) provides a standardized pipeline to evaluate imputation performance and offers a systematic way to adapt forecasting algorithms for imputation. It involved over 34,000 experiments, 28 algorithms, and 8 datasets.
@@ 6.4 Online Imputation for Streaming Data
Online data imputation refers to the process of estimating and filling in missing values in a continuous, streaming data environment where new data arrives sequentially over time, and the imputation model must update its estimates incrementally without reprocessing the entire dataset. Compared to batch or offline imputation, online imputation should adapt and learn from new incomplete data dynamically. That means online imputation enables real-time decision-making, monitoring, and analytics in dynamic environments. To be more specific, suppose we have one observation $\mathbf{x}_{t}\in\mathbb{R}^{m}$ at time point $t$, of which the missing entries are indicated by $\bm{\omega}_{t}\in\{0,1\}^{m}$. We need to construct a model $\mathcal{M}_{t}$ to impute the missing values of $\mathbf{x}_{t}$ using $\{\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{t-1}\}$ quickly, where $\mathcal{M}_{t}$ evolves with $t$ since the distribution of the complete data may change over time, usually slowly. Online missing data imputation finds critical applications in domains such as IoT and sensor networks (Deng et al., 2022), chemical process monitoring, financial trading (Nishanth and Ravi, 2013), and patient health monitoring.
A straightforward approach to adapting the aforementioned offline imputation methods for an online setting is to update the model incrementally, for instance via gradient descent, as new data arrives. Such an approach, however, requires careful design to be effective. Over the past decades, several online imputation algorithms have been proposed. For instance, Brand (2003) developed a fast online SVD revision method for missing data imputation, which was applied to collaborative filtering, allowing users to asynchronously join, add ratings, add movies, revise ratings, get recommendations, and delete themselves from the model. Balzano et al. (2010) provided an algorithm of online subspace tracking on incomplete data, which can be used for online missing data imputation. Ravi and Krishna (2014) developed an auto associative neural network method for online missing data imputation. Guo (2015) developed an online LRMC method based on low-rank factorization. Devooght et al. (2015) developed an algorithm of dynamic matrix factorization for collaborative filtering. Fan and Udell (2019a) proposed an online KFMC method to recover high rank matrices online. Zhao et al. (2022) provided a Gaussian Copula-based online imputation algorithm for data of mixed types, including ordinal, boolean, and continuous variables. Fan (2022b) presented a dynamic nonlinear matrix completion for online data imputation based on fast eigenvalue decompositions of the kernel matrices constructed by a sliding window of the data stream. Liu et al. (2023b) proposed an online imputation algorithm that is able to handle binary, count and continuous variables simultaneously. Zhan et al. (2025b) developed a mask asymmetric transformer GAN for online imputation in edge computing.
@@ 6.5 Categorical Data Imputation
Categorical data imputation (Van Buuren and van Rijckevorsel, 1992; Chen and Åstebro, 2003; Cranmer and Gill, 2013) is fundamentally more intricate than numerical data imputation due to the following reasons. First, unlike numerical variables, categorical data lack intrinsic ordering or meaningful distance, so mean-based or regression-style imputations are invalid and even model-based solutions must first convert categories to dummies whose continuous outputs then have to be rounded while preserving mutual exclusivity, a step that introduces bias. Second, missingness mechanisms are harder to specify or test for categorical data, since the unobserved truth is discrete. Lastly, categorical imputation is harder to evaluate: a single mis-classification counts as a full error, so we must rely on high-variance metrics like hit-rate, F1-score or Cohen’s $\kappa$ coefficient instead of the stable RMSE/MAE used for numerical data.
Classical methods for categorical data imputation include kNN with Gower or Hamming distance (Schwender, 2012; Faisal and Tutz, 2022), CART (classification and regression trees (Breiman et al., 2017), MICE (van Buuren and Groothuis-Oudshoorn, 2011), and graphical models (Geng et al., 2003). Akande et al. (2017) and Memon et al. (2023) conducted empirical comparisons of several classical methods for categorical data. Nishanth and Ravi (2016) proposed a probabilistic neural network for categorical data imputation. The method outperformed classical methods such as kNN. Note that by converting all categorical variables into numerical form using techniques like one-hot encoding, we can perform numerical imputation methods, especially those based on deep neural networks. It is also possible to use LLMs to impute categorical data, as discussed in Section 5.7.
@@ 6.6 Multimodal Data Imputation
Multimodal data, characterized by the integration of diverse data types such as images, text, tabular measurements, and graphs, is ubiquitous in modern artificial intelligence applications. Examples include electronic health records (EHRs) combining clinical measurements (tabular), medical notes (text), and X-ray images (vision); social media profiles with user demographics (tabular), posts (text), and profile pictures (vision); and product catalogs with specifications (tabular), descriptions (text), and customer interaction graphs. Learning from multimodal data is more effective than from a single modality (Ngiam et al., 2011; Baltrušaitis et al., 2018). However, missing data in multimodal settings presents a uniquely complex challenge. The missingness can occur not only within individual features but across entire modalities for certain data instances. For instance, one patient in an EHR may have complete lab tests but no associated medical image, while another may have the reverse.
@@
Figure 10 presents a toy example of multimodal data imputation. As can be seen, multimodal data imputation is significantly more intricate than unimodal imputation. The former may require an integration of multiple machine learning models or techniques, e.g., a combination of a convolutional neural network and a recurrent neural network for handling image and text simultaneously. Zhan et al. (2025a) provided a literature review on incomplete multimodal learning. It categorizes the related methods into two groups: internal information-based methods and external information-based methods. Since our paper focuses on missing data imputation, we here detail the internal information-based methods, particularly how to impute the multimodal data, without using any external information.
@@
Figure 10: An intuitive example of multimodal data imputation (three modalities, ten subjects). Each cell represents a numerical (or categorical) value, an image, or a text. The content in every cell marked by the red cross is missing.
@@
Tran et al. (2017) presented a multimodal imputation method based on a cascaded residual autoencoder. Shang et al. (2017) proposed combining GANs and AEs for multimodal imputation, where the AEs for different modalities share a common representation space to exploit cross-modal relationships. Chen et al. (2023) proposed a method that disentangles brain images into inter-modality relevant and intra-modality specific features, and uses knowledge distillation to impute missing modality representations, enabling multimodal Alzheimer’s disease diagnosis using only single-modality data at inference. Cohen Kalafut et al. (2023) developed a machine learning model based on joint VAEs for single-cell multimodal data imputation and embedding. Similarly, Tang et al. (2024) combined contrastive learning within modality-specific AEs for single-cell data imputation. More work about multi-omics data imputation can be found in (Du et al., 2022; Chandrashekar et al., 2023; Wang et al., 2024b; Jiang et al., ; Sadia and Cheng, 2025). Poudel et al. (2025) proposed an imputation method for multimodal federated learning and evaluated it on three medical datasets containing text and image modalities. Hassanzadeh et al. (2025) provided a cycle-GAN imputation method for Alzheimer’s disease diagnosis based on two modalities. Zhang et al. (2025d) applied diffusion models to federated multimodal graph completion, utilizing image, text, and graph information simultaneously. There is additional work on handling multimodal data, primarily focused on heuristic-based applications such as medical diagnosis, which we will not elaborate on in this paper.
@@
The principle followed by most multimodal data imputation methods is that different modalities are correlated or share some common information; otherwise, it is impossible to recover one modality from others. Specifically, let $\mathbf{x}^{(j)}\in\mathcal{X}_{j}$ be the data of a subject or sample in modality $j$; $\mathbf{x}^{(j)}$ could be a vector, an image, a text, a graph, or a time series, etc, where $j=1,2,\ldots,M$ and $M$ is the number of modalities. Let $\mathbf{x}=(\mathbf{x}^{(1)},\mathbf{x}^{(2)},\ldots,\mathbf{x}^{(M)})$ and let $\mathcal{D}=\{\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{N}\}$ be the whole dataset of the multimodal data consisting of $N$ samples. Suppose $\mathcal{D}$ is incomplete, where an $\mathbf{x}_{i}^{(j)}$ may have some missing elements or is completely missing. Most deep learning-based missing data imputation methods we introduced in Section 5.6 can be extended to impute $\mathcal{D}$. For instance, we consider an autoencoder-type imputation strategy. Let $f:\mathcal{X}_{1}\times\mathcal{X}_{2}\times\cdots\mathcal{X}_{M}\rightarrow\mathbb{R}^{d}$ be an encoder with $M$ different types of input and $g:\mathbb{R}^{d}\rightarrow\mathcal{X}_{1}\times\mathcal{X}_{2}\times\cdots\mathcal{X}_{M}$ be a decoder. Then we have the following framework for multimodal data imputation:
$\mathop{\text{minimize}}_{f,g}\frac{1}{N}\sum_{i=1}^{N}\ell(\mathbf{x}_{i},g(f(\mathbf{x}_{i})),\bm{\omega}_{i})$ (34)
@@
where $\ell(\mathbf{x},\hat{\mathbf{x}},\bm{\omega})$ denotes a function to compare $\mathbf{x}$ with its reconstruction $\hat{\mathbf{x}}$ only at the observed locations indicated by $\bm{\omega}$. It should be pointed out that $\ell$ is not a simple loss function such as the square loss; it should be able to compare two images, two texts, or even two graphs; $\ell$ itself could be a multimodal neural network, pretrained, or jointly optimized with $f$ and $g$. This idea can be adapted to other approaches, such as GAN and diffusion models, to impute multimodal data.
@@ 7 Supervised and Unsupervised Learning on Incomplete Data
Missing data imputation often has a significant impact on downstream tasks such as classification and clustering, especially when the missing rates are high (Ghahramani and Jordan, 1995; Saar-Tsechansky and Provost, 2007; García-Laencina et al., 2009a). Performing downstream task-agnostic imputation and then conducting the downstream task may not provide satisfactory performance. The conventional approach of employing task-agnostic imputation strategies followed by separate downstream analysis frequently yields suboptimal performance outcomes. To address the limitation of this two-stage approach, one may need to ensure that the assumption on the data distribution or structure is the same as that in the downstream task. Moreover, a few researchers proposed end-to-end approaches that integrate missing data imputation with the downstream tasks.

### v2

@@ 6.2 Link Prediction and Graph Completion
Graph completion, which is more general than link prediction, aims to enrich the entire graph by adding not only edges but also missing nodes, attributes, or labels so that the resulting graph is more complete and useful for downstream tasks. A general graph is formulated as $G=(V,E,\mathbf{X})$, where $\mathbf{X}$ denotes the matrix of node attributes. Graph completion may recover the missing edges as well as the missing values in $\mathbf{X}$, or predict where a new node should be inserted into the graph (wu2023overview). Within the tasks of graph completion, knowledge graph completion (KGC) is best known. KGC is the task of predicting missing facts in a knowledge graph, typically by inferring new relationships between entities to answer queries like (subject, predicate, ?) or (?, predicate, object). There have been many review papers on KGC (9220143; SHEN2022109597; zamini2022review; cai2022temporal; 10.1007/s00521-023-09286-2; LI2025100809). In (LI2025100809), KGC methods are organized into embedding-based methods, path-based methods, neural network-based methods, and LLM-based methods (wang2022simkgc; wei2024kicgpt; xu2024multi; li2024contextualization; zhang2024making; yao2025exploring).
@@ 6.3 Time Series Imputation
The task of time series imputation is to recover the missing values in a single multivariate time series (e.g., a matrix $\mathbf{X}\in\mathbb{R}^{d\times T}$) or a union of independent multivariate time series (e.g., a union of matrices $\mathbf{X}_{i}\in\mathbb{R}^{d\times T_{i}}$, $i=1,\ldots,K$). Compared to the standard missing data imputation, time series data imputation should consider the auto-correlation, trend/seasonality, and lagged cross-correlation. Time-series imputation keeps critical real-time systems (power grids, ICU monitors, high-frequency trading) from blinding when sensors drop out, preventing cascading failures or false alarms. It also unlocks the full value of expensive historical records—satellite climate data, industrial IoT (internet of things) logs, retail POS streams—so forecasting, digital twins and compliance audits can run on complete, audit-ready data instead of discarding gap-ridden sequences. Recent advances in time series imputation are based on recurrent neural networks (cao2018brits) and diffusion models (tashiro2021csdi). There have been a few review papers on time series imputation (pratama2016review; weerakody2021review; ribeiro2022missing; zainuddin2022time; fang2020time; wang2024deep; qian2025deep). In the past ten years, deep learning methods, especially deep generative models, have shown impressive performance in time-series imputation. fang2020time wrote a survey paper of deep learning-based imputation methods for time series (luo2018multivariate; gupta2020time; park2023long). kazijevs2023deep focused on time series health data and provided a review with numerical comparisons for a few deep learning imputation methods. jin2024survey reviewed graph neural network-based methods for time series imputation. du2024tsi provides a standardized pipeline to evaluate imputation performance and offers a systematic way to adapt forecasting algorithms for imputation. It involved over 34,000 experiments, 28 algorithms, and 8 datasets.
@@ 6.4 Online Imputation for Streaming Data
Online data imputation refers to the process of estimating and filling in missing values in a continuous, streaming data environment where new data arrives sequentially over time, and the imputation model must update its estimates incrementally without reprocessing the entire dataset. Compared to batch or offline imputation, online imputation should adapt and learn from new incomplete data dynamically. That means online imputation enables real-time decision-making, monitoring, and analytics in dynamic environments. To be more specific, suppose we have one observation $\mathbf{x}_{t}\in\mathbb{R}^{m}$ at time point $t$, of which the missing entries are indicated by $\bm{\omega}_{t}\in\{0,1\}^{m}$. We need to construct a model $\mathcal{M}_{t}$ to impute the missing values of $\mathbf{x}_{t}$ using $\{\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{t-1}\}$ quickly, where $\mathcal{M}_{t}$ evolves with $t$ since the distribution of the complete data may change over time, usually slowly. Online missing data imputation finds critical applications in domains such as IoT and sensor networks (deng2022online), chemical process monitoring, financial trading (nishanth2013computational), and patient health monitoring.
A straightforward approach to adapting the aforementioned offline imputation methods for an online setting is to update the model incrementally, for instance via gradient descent, as new data arrives. Such an approach, however, requires careful design to be effective. Over the past decades, several online imputation algorithms have been proposed. For instance, brand2003fast developed a fast online SVD revision method for missing data imputation, which was applied to collaborative filtering, allowing users to asynchronously join, add ratings, add movies, revise ratings, get recommendations, and delete themselves from the model. MC_GROUSE_2010 provided an algorithm of online subspace tracking on incomplete data, which can be used for online missing data imputation. ravi2014new developed an auto-associative neural network method for online missing data imputation. guo2015online developed an online LRMC method based on low-rank factorization. devooght2015dynamic developed an algorithm of dynamic matrix factorization for collaborative filtering. Fan_2019_CVPR proposed an online KFMC method to recover high rank matrices online. zhao2022online provided a Gaussian Copula-based online imputation algorithm for data of mixed types, including ordinal, boolean, and continuous variables. fan2022dynamic presented a dynamic nonlinear matrix completion for online data imputation based on fast eigenvalue decompositions of the kernel matrices constructed by a sliding window of the data stream. LIU2023107822 proposed an online imputation algorithm that is able to handle binary, count and continuous variables simultaneously. zhan2025online developed a mask asymmetric transformer GAN for online imputation in edge computing.
@@ 6.5 Categorical Data Imputation
Categorical data imputation (van1992imputation; chen2003deal; cranmer2013we) is fundamentally more intricate than numerical data imputation due to the following reasons. First, unlike numerical variables, categorical data lack intrinsic ordering or meaningful distance, so mean-based or regression-style imputations are invalid and even model-based solutions must first convert categories to dummies whose continuous outputs then have to be rounded while preserving mutual exclusivity, a step that introduces bias. Second, missingness mechanisms are harder to specify or test for categorical data, since the unobserved truth is discrete. Lastly, categorical imputation is harder to evaluate: a single misclassification counts as a full error, so we must rely on high-variance metrics, like hit-rate, F1-score, or Cohen’s $\kappa$ coefficient, instead of the stable RMSE/MAE used for numerical data.
Classical methods for categorical data imputation include kNN with Gower or Hamming distance (schwender2012imputing; faisal2022nearest), CART (classification and regression trees (breiman2017classification), MICE (MICE), and graphical models (geng2003bayesian). akande2017empirical and memon2023comparison conducted empirical comparisons of several classical methods for categorical data. nishanth2016probabilistic proposed a probabilistic neural network for categorical data imputation. The method outperformed classical methods such as kNN. Note that by converting all categorical variables into numerical form using techniques like one-hot encoding, we can perform numerical imputation methods, especially those based on deep neural networks. It is also possible to use LLMs to impute categorical data, as discussed in Section 5.7.
@@ 6.6 Multimodal Data Imputation
Multimodal data, characterized by the integration of diverse data types such as images, text, tabular measurements, and graphs, is ubiquitous in modern artificial intelligence applications. Examples include electronic health records (EHRs) combining clinical measurements (tabular), medical notes (text), and X-ray images (vision); social media profiles with user demographics (tabular), posts (text), and profile pictures (vision); and product catalogs with specifications (tabular), descriptions (text), and customer interaction graphs. Learning from multimodal data is more effective than from a single modality (ngiam2011multimodal; baltruvsaitis2018multimodal). However, missing data in multimodal settings presents a uniquely complex challenge. The missingness can occur not only within individual features but across entire modalities for certain data instances. For instance, one patient in an EHR may have complete lab tests but no associated medical image, while another may have the reverse.
@@
Figure 10 presents a toy example of multimodal data imputation. As can be seen, multimodal data imputation is significantly more intricate than unimodal imputation. The former may require an integration of multiple machine learning models or techniques, e.g., a combination of a convolutional neural network and a recurrent neural network for handling image and text simultaneously. zhan2025systematic provided a literature review on incomplete multimodal learning. It categorizes the related methods into two groups: internal information-based methods and external information-based methods. Since our paper focuses on missing data imputation, we here detail the internal information-based methods, particularly how to impute the multimodal data, without using any external information.
@@
Figure 10: An intuitive example of multimodal data imputation (three modalities, ten subjects). Each cell represents a numerical (or categorical) value, an image, or text. The content in every cell marked by the red cross is missing.
@@
tran2017missing presented a multimodal imputation method based on a cascaded residual autoencoder. shang2017vigan proposed combining GANs and AEs for multimodal imputation, where the AEs for different modalities share a common representation space to exploit cross-modal relationships. chen2023disentangle proposed a method that disentangles brain images into inter-modality relevant and intra-modality specific features, and uses knowledge distillation to impute missing modality representations, enabling multimodal Alzheimer’s disease diagnosis using only single-modality data at inference. cohen2023joint developed a machine learning model based on joint VAEs for single-cell multimodal data imputation and embedding. Similarly, tang2024modal combined contrastive learning within modality-specific AEs for single-cell data imputation. More work about multi-omics data imputation can be found in (du2022robust; chandrashekar2023deepgami; wang2024inclust+; jiang2024xmint; sadia2025depmicrodiff). poudel2025multimodal proposed an imputation method for multimodal federated learning and evaluated it on three medical datasets containing text and image modalities. hassanzadeh2025generative provided a cycle-GAN imputation method for Alzheimer’s disease diagnosis based on two modalities. zhang2025hyper applied diffusion models to federated multimodal graph completion, utilizing image, text, and graph information simultaneously. There is additional work on handling multimodal data, primarily focused on heuristic-based applications such as medical diagnosis, which we will not elaborate on in this paper.
@@
The principle followed by most multimodal data imputation methods is that different modalities are correlated or share some common information; otherwise, it is impossible to recover one modality from others. Specifically, let $\mathbf{x}^{(j)}\in\mathcal{X}_{j}$ be the data of a subject or sample in modality $j$; $\mathbf{x}^{(j)}$ could be a vector, an image, a text, a graph, or a time series, etc, where $j=1,2,\ldots,M$ and $M$ is the number of modalities. Let $\mathbf{x}=(\mathbf{x}^{(1)},\mathbf{x}^{(2)},\ldots,\mathbf{x}^{(M)})$ and let $\mathcal{D}=\{\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{N}\}$ be the whole dataset of the multimodal data consisting of $N$ samples. Suppose $\mathcal{D}$ is incomplete, where an $\mathbf{x}_{i}^{(j)}$ may have some missing elements or is completely missing. Most deep learning-based missing data imputation methods we introduced in Section 5.6 can be extended to impute $\mathcal{D}$. For instance, here we consider an autoencoder-type imputation strategy. Let $f:\mathcal{X}_{1}\times\mathcal{X}_{2}\times\cdots\mathcal{X}_{M}\rightarrow\mathbb{R}^{d}$ be an encoder with $M$ different types of input and $g:\mathbb{R}^{d}\rightarrow\mathcal{X}_{1}\times\mathcal{X}_{2}\times\cdots\mathcal{X}_{M}$ be a decoder. Then we have the following framework for multimodal data imputation:
$\mathop{\text{minimize}}_{f,g}~\frac{1}{N}\sum_{i=1}^{N}\ell(\mathbf{x}_{i},g(f(\mathbf{x}_{i}))|\bm{\omega}_{i})$ (45)
@@
where $\ell(\mathbf{x},\hat{\mathbf{x}}|\bm{\omega})$ denotes a function to compare $\mathbf{x}$ with its reconstruction $\hat{\mathbf{x}}$ only at the observed locations indicated by $\bm{\omega}$. It should be pointed out that $\ell$ is not a simple loss function such as the square loss; it should be able to compare two images, two texts, or even two graphs; $\ell$ itself could be a multimodal neural network, pretrained, or jointly optimized with $f$ and $g$. This idea can be adapted to other approaches, such as GAN and diffusion models, to impute multimodal data.
@@ 7 Supervised and Unsupervised Learning on Incomplete Data
Missing data imputation often has a significant impact on downstream tasks such as classification and clustering, especially when the missing rates are high (ghahramani1995learning; saar2007handling; EM). Performing downstream task-agnostic imputation and then conducting the downstream task may not provide satisfactory performance. The conventional approach of employing task-agnostic imputation strategies followed by separate downstream analysis frequently yields suboptimal performance outcomes. To address the limitation of this two-stage approach, one may need to ensure that the assumption on the data distribution or structure is the same as that in the downstream task. Moreover, a few researchers proposed end-to-end approaches that integrate missing data imputation with downstream tasks.